OpenStudy (taepoohwae):

Can someone please show me how to solve this problem y^2-16 step by step by factoring a difference of squares so that I have a better understanding? Betore asking, no I do not know what to do.

OpenStudy (anonymous):

are we solving for 0?

OpenStudy (anonymous):

Just pretend y^2-16=0 Now you factor the problem, getting (y-4)(y+4)=0 Now take each of these separately to equal 0 y-4=0 and y+4=0 Now solve for y in each equation. You get y=4 and -4 Because of the law of something(I forgot what it is...But trust me, it exists), you know know that (y-4) and (y+4) are both factors of y^2-16

OpenStudy (anonymous):

Or you could add 16 to both sides, giving you y^2=16 Now you find the square root of both sides and voila! y=4 and/or -4

OpenStudy (taepoohwae):

It's factoring... I wrote the wrong problem also. It should be y^2+72

OpenStudy (anonymous):

Then subract 72 from both sides of y^2+72=0 and find the square root of both sides y^2=-72 Because you cannot find the square root of a negative number, you make the problem \[\sqrt{y^2}=\sqrt{(-1)(72)}\]

OpenStudy (anonymous):

The square root of negative 1 is also known as i. So now you have \[\sqrt{y^2}=i \sqrt{72}\]

OpenStudy (anonymous):

Now you solve for the square root 72

OpenStudy (anonymous):

If I just confused you terribly, let me know.

OpenStudy (taepoohwae):

Yes, because from my understanding the answer ends (a+b)(a-b)